# Agentic AI Software Engineer: Programming with Trust

### v2

@@ Abstract.
Large Language Models (LLMs) have shown surprising proficiency in generating code snippets, promising to automate large parts of software engineering via artificial intelligence (AI). We argue that successfully deploying AI software engineers requires a level of trust equal to or even greater than the trust established by human-driven software engineering practices. The recent trend toward LLM agents offers a path toward integrating the power of LLMs to create new code with the power of analysis tools to increase trust in the code. This opinion piece comments on whether LLM agents could dominate software engineering workflows in the future and whether the focus of programming will shift from programming at scale to programming with trust.
1. Are AI Software Engineers Deployed?
Software engineering is undergoing a significant phase of greater automation owing to the emergence of Large Language Models (LLMs) for code. It is easy to generate almost correct code by feeding in natural language requirements to LLMs. Such automated code generation creates public excitement and shows the promise of automatic programming. Given the recent progress, some are already declaring victory and argue that AI has “solved” the software development problem.
@@ 1. Are AI Software Engineers Deployed?
While genAI-enabled code generation and code completion are now prevalent in common IDEs, fully automated AI software engineering is not yet widely deployed in industrial practice. What is holding people back from adopting generative AI? A recent blog post (Kohler, 2025) by the behavioral scientist and future of work advocate Lindsay Kohler points out that the key barrier to genAI adoption is largely trust. People are asking if they can trust AI and if they can demonstrate trustworthiness to stakeholders. In the domain of coding, the concern is thus not about the management of the organization not accepting automatic programming, but it is about automatic programming not being accepted by developers, due to lack of trust.
@@
This brings us to the following question: What is the place of automatically generated code in future development workflows? If we can figure out how automatically generated code and manually written code can co-exist, this may give us a pathway of greater deployment of LLMs for software engineering! Starting from early programs of just a few lines written in high-level languages in the 1950s, size of programs has increased greatly to hundreds of millions of lines of code. For the past fifty years, there has been clear interest towards programming in the large, where distributed software development by different teams is carried out across different modules in a software project. For programming in the large, the focus is often on software project management and cleanly partitioning work among developer teams. As a result, for the past fifty years, a lot of attention in software engineering has been on managing the complexities of scale in large software projects. With the arrival of AI software engineers, programming is less about code writing and more about code assembling. Importantly, a lot of the assembled code will be generated automatically or come from other potentially untrusted sources. This development turns the attention towards the trust that can be placed on such component assemblies.
It is the opinion of these authors that increased automation in programming will make programming with trust at least as important as, if not more important than, programming at scale.
2. Agentic Capability
@@
As LLMs alone cannot establish sufficient trust by developers to produce correct code, we envision a role for LLM agents in genAI-based software engineering. What is an LLM agent? How does the use of LLM agents differ from prompt engineering for generating code? We highlight the following aspects about LLM agents for software.111See also the following for a discussion https://www.anthropic.com/research/building-effective-agents
LLMs as back-ends: An LLM agent is a wrapper around LLMs, using LLMs as back-end computation engines. Agents can interact with multiple LLMs in the back-end.
Interaction with software tools: Agents interact with different tools to achieve a given task. In software engineering, the agentic capability often comes from invoking software testing and analysis tools. Other, more generic tools, can also be used, such as using a shell tool to navigate files in a code base or to invoke arbitrary command-line tools. Appropriate use of testing and analysis tools can enhance trust of the developer in the results of the LLM agent.
@@
Autonomy: An agent invokes tools in an autonomous manner. That is, the agent is not a deterministic algorithm, but rather creates a nondeterministic work-plan with significant autonomy. Deciding on the degree of autonomy is part and parcel of designing an LLM agent.
@@
Recently, several software LLM agents have been proposed, starting with the announcement of the Devin AI software engineer from Cognition Labs (Wu, 2024). Devin can solve natural language tasks (called issues) involving bug-fixes and feature-additions. The approach combines a back-end LLM with access to standard developer tools, such as a shell, a code editor, and a web browser. Such tools will be employed autonomously to let the AI software engineer mimic human software engineers. In parallel with the announcement of Devin, several research groups proposed their own LLM agents for software engineering, including RepairAgent (Bouzenia et al., 2025), AutoCodeRover (Zhang et al., 2024), SWE-agent (Yang et al., 2024), and Agentless (Xia et al., 2025).
RepairAgent (Bouzenia et al., 2025) focuses on fixing bugs and guides the LLM agent by defining a finite-state machine that outlines the typical steps followed by a developer. In each state, only a subset of all tools, such as code editing, code search, and testing, is available to the agent. RepairAgent cannot solve natural language issues, and thus cannot behave like a human software engineer. Instead it focuses on conventional program repair where the goal is to make failing tests pass.
AutoCodeRover, a spinoff acquired by SonarSource, can solve natural language issues requiring bug fixing or feature addition. It was reported to have higher efficacy (Zhang et al., 2024) than Devin, owing largely to its attention to extracting the software intent. To this end, the AI software engineer is equipped with program analysis techniques, such as autonomous code search (on an abstract syntax tree) and spectrum-based fault localization. Such intent extraction for software engineering tasks can also allow the agent to provide explanations of fixes.
SWE-agent (Yang et al., 2024) follows a philosophy similar to Devin, by making file navigation tools and interfaces available to an AI software engineer. Thus, it does not employ any program analysis and does not work on program representations.
Last but not least, the Agentless approach (Xia et al., 2025) advocates the use of careful prompt engineering to LLMs for fault localization, repair, and patch validation. This is done without allowing autonomy to an agent operating on top of an LLM. While there is no autonomy, Agentless is nevertheless an AI software engineer conducting software maintenance and evolution.
@@ 3. Establishing Trust
There exist numerous examples of LLM-generated code that suffers from bugs and security vulnerabilities. Obviously, human-written code may also suffer from these problems. What makes us trust human-written, but not LLM-generated code? Part of the reason is the perceived capability of “passing the blame”. If a human developer is involved, there is the promise of getting feedback from the developer as needed. Of course, this does not always hold, e.g., if the developer eventually leaves an organization. Nevertheless, accepting a code commit from a developer often partially depends on the reputation of the developer in the organization. For AI software engineers, until they become commonplace, there is a greater need to engender trust by applying quality assurance techniques like testing, static analysis, and formal verification. So how can AI software engineers alleviate the issue of trust in automatically generated code? There can be varying levels of trust that the AI software engineer can establish.
Testing for Consistency
At a technical level, some initial degree of trust can be ensured by retrofitting testing into AI software engineering workflows. Thus, in the process of code generation, additional artifacts, such as tests, can be generated as well (Ryan et al., 2024). For the generated tests, not only the test inputs, but also the expected outputs, i.e., the oracles, of the tests will be generated from natural language specifications. These tests can be cross-checked with the bug fixes or feature additions generated by the AI software engineer, thereby yielding greater trust in the program modifications.
Specification Inference
A more conceptual mechanism to establish trust would be to infer the code intent from the initial, possibly buggy program. The system-level intent of what a large software system is supposed to do can often be crisply captured by a detailed natural language prompt. What is missing is the intent of the different code units, such as functions or methods. An AI software engineer could be geared towards such specification inference, navigating the code-base via code search, and trying to infer intended pre/post-conditions (Ruan et al., 2025). Such explicit specification inference can enable the program modifications made by AI software engineers to be accompanied by well-explained justifications, increasing trust. The extracted specifications can also be used to cross-check the oracles of any generated tests.
@@
An enhanced degree of trust can come from formal proofs. A promising paradigm in this regard is automated, proof-oriented programming (Chakraborty et al., 2025), where the LLMs are used to generate code together with a proof that can be automatically checked with an external tool. While historically, proving program correctness has been very difficult due to the high cognitive burden on programmers, LLMs can help by generating the code together with the necessary assertions (pre/post-conditions, loop invariants, and so on) in a verifiable language such as F*, Dafny, or Verus. Such programs can then be automatically verified, providing greater trust than testing alone.
Guardrails for Increased Security
Trust can also be ensured through the use of guardrails.222See https://hub.guardrailsai.com/ for example. These can serve as sanitization mechanisms, filtering malicious inputs before they reach the LLM and validating the generated code. Guardrails can be used to protect against the following threats: (i) Prompt injections: Natural language prompts that jailbreak the LLM, convincing it to disregard its safety policies and to perform malicious activities, like generating malicious or vulnerable code. (ii) Malicious code: Even if the natural language part of the prompt is benign, input code snippets may be malicious, forcing the LLM to generate (or modify) malicious code. (iii) Vulnerable code – Even without any malicious intent, the input code may still be vulnerable, forcing the LLM to modify vulnerable code. Static checkers can be used for guardrailing such malicious behavior.
Open Challenges
While we see establishing trust as the prime challenge for widely deploying LLM agents in software engineering, there are some related challenges that also need to be addressed: (i) Communication: The explanations given by the AI software engineer must be communicated properly without imposing high cognitive load. Checking the results of AI-generated code currently involves a significant human effort. (ii) Diverse teams: Using AI Software Engineer excessively may reduce diversity of thought333See https://medium.com/@LuisMizutani/ai-for-software-development-can-we-actually-measure-its-impact-31af45dccdab in a development team. Thus, co-operative intelligence across AI-AI and AI-human combinations remains a topic of discourse. (iii) Costs: Even though the automation provided by AI promises to reduce costs, an AI software engineer still needs to be cost-sensitive. Program analysis tools invoked by LLM agents can help make only relevant queries to LLMs (Zhang et al., 2024). (iv) Datasets: Last but not the least, the community needs datasets. The latest dataset announced in this regard is the SWE-Lancer (Miserendino et al., 2025). We need datasets for software engineering tasks which support comprehensive patch validation (possibly by supporting test generation), and, ideally, cover more capabilities of the AI software engineer than patching, as we discuss in the following. .
@@ 4. Outlook
With LLM agents performing some of the heavy lifting in software maintenance and software quality control, it may be useful to think about how they will fit into future development workflows. Will there be separate agents for coding, testing, debugging, repair, and various software engineering tasks? The more agents there are, the more the maintainability of these agent may become an issue! Instead of building hyper-specialized agents for individual software engineering tasks, it could be worthwhile to craft and maintain a unified software engineering agent (USE-agent) that can be configured to deliver different capabilities. Such an agent of the future may conduct myriad tasks, including code generation, test generation, bug fixing, and feature addition. Software engineering can then be conducted by clients that rely on the unified agent, with human developers vetting the results of the agent wherever needed. With newer models like o1 and DeepSeek R1 being developed, some of the current agentic capabilities can be co-opted by LLMs in the future, and hence it is useful to think about more ambitious, unified agents.
@@ Acknowledgements.
- Xia et al. (2025) Chunqiu Steven Xia, Yinlin Deng, Soren Dunn, and Lingming Zhang. 2025. Agentless: Demystifying LLM-based Software Engineering Agents. In 33rd ACM SIGSOFT International Symposium on the Foundations of Software Engineering (FSE).

### v3

# Agentic AI Software Engineers: Programming with Trust

@@ Abstract.
Large Language Models (LLMs) have shown surprising proficiency in generating code snippets, promising to automate large parts of software engineering via artificial intelligence (AI). We argue that successfully deploying AI software engineers requires a level of trust equal to, or even greater than, the trust established by human-driven software engineering practices. The recent trend toward LLM agents offers a path toward integrating the power of LLMs to create new code with the power of analysis tools to increase trust in the code. This opinion piece comments on how LLM agents present AI software engineering workflows of the future, and whether the focus of programming will shift from scale to trust.
1. A Key Barrier for AI Software Engineers
@@
Software engineering is undergoing a disruptive phase of greater automation owing to the emergence of Large Language Models (LLMs) that generate and edit code. This progress creates public excitement about AI software engineers, which promise to largely automate many core software development tasks, potentially saving tremendous costs (Miserendino et al., 2025). While AI-enabled code generation and code editing are now prevalent in integrated development environments (IDEs), fully automated AI software engineers are not yet widely deployed in industrial practice. What is holding people back from adopting AI software engineers? A recent blog post by the behavioral scientist and future-of-work advocate Lindsay Kohler points out that the key barrier to AI adoption is trust (Kohler, 2025). Users are wondering if they can trust AI, and how they can demonstrate trustworthiness to stakeholders. In the domain of software engineering, the concern is thus not about the management of an organization not accepting AI software engineers, but it is about developers not trusting their new AI companions.
@@
This brings us to the following question: What is the place of AI software engineers in future development workflows? If we can figure out how automatically generated and manually written software can co-exist, this may give us a pathway of greater deployment of AI in software engineering! Starting from early programs of just a few lines written in high-level languages in the 1950s, the size of programs has increased greatly to hundreds of millions of lines of code. For the past fifty years, there has been a steady interest towards programming in the large, where distributed software development by different teams is carried out across different modules of a software project. With the increased use of AI code generation, we believe that the emphasis will be not only on programming at scale, but increasingly on programming with trust.
2. Technical and Human Trust
| Aspect | Trust Factor |
| Technical | Correctness: Does the code produce the expected results? |
| Security: Is the code free of vulnerabilities and safe to deploy? | |
| Performance: Is the code efficient and scalable? | |
| Maintainability: Is the code easy to read, refactor, and extend? | |
| Compliance: Does the code adhere to language/framework best practices? | |
| Testing: Are there sufficient and meaningful test cases? | |
| Dependencies: Are dependencies safe, current, and widely trusted? | |
| Human | Explainability and Transparency: Can the AI justify the chosen solution? |
| Bias and Ethics: Does the code reflect unbiased, ethical decisions? | |
| Familiarity: Are the code style or algorithms familiar to the developer? | |
| Over-Reliance: Is there concern the user might use the code blindly? | |
| Experience Mismatch: Does the code align with the developer’s expertise level? | |
| Collaboration: Does the AI react well to developer feedback? | |
| Team Practice: Does the AI align with the team’s development workflow, conventions, and review culture? |
Table 1 illustrates how developers perceive trust in AI-generated code, both from a technical and a human angle. Technical trust in AI-generated code stems from measurable attributes that determine code quality and reliability. Developers need assurance that the code is correct, secure, and performs efficiently under expected conditions. Maintainability and standards compliance are crucial for long-term usability, while thorough testing validates that the code functions as intended. Additionally, trust improves when dependencies are well-managed and sourced from reputable libraries.
Human trust involves psychological and social dynamics that influence how developers perceive and accept AI-generated solutions. A key factor is explainability and transparency: When an AI system can clarify its reasoning and design choices, developers are more likely to trust and adopt its outputs. Developers are also more likely to trust code that aligns with ethical values, follows familiar patterns and team practices, and matches the developers level of expertise. An important risk that could reduce trust is over-reliance on AI-provided code without properly understanding it. Furthermore, trust is strengthened when AI systems collaborate effectively and adapt to feedback provided by developers.
3. Software Engineering Agents
@@
As LLMs alone cannot inspire sufficient trust, we see LLM agents for software engineering as a promising way of creating trustworthy AI software engineers. What is an LLM agent for software, and how does it differ from prompt engineering? We highlight three aspects of such agents:111See also the following for a discussion https://www.anthropic.com/research/building-effective-agents
LLMs as back-ends: An agent is a program that leverages one or more LLMs as back-end computation and decision engines.
Interaction with software tools: An agent interacts with different tools to achieve a given task. In software engineering, such tools resemble those commonly used by human software engineers, e.g., file navigation, code editing, executing test suites, and invoking program analysis tools. Appropriate use of these tools is key for enhancing trust of developers in the results of the LLM agent.
@@
Autonomy: An agent invokes tools in an autonomous manner. That is, the agent does not follow a deterministic algorithm, but rather creates a nondeterministic work-plan with significant autonomy.
@@
Recently, several software engineering agents have been proposed, starting with the announcement of the Devin AI software engineer from Cognition Labs (Wu, 2024). Devin can solve natural language tasks (called issues), such as bug fixes and feature additions. It combines a back-end LLM with access to standard developer tools, such as a shell, a code editor, and a web browser. The agent employs such tools autonomously to let the AI software engineer mimic human practices. In parallel with the announcement of Devin, several research groups proposed their own LLM agents for software engineering, including RepairAgent (Bouzenia et al., 2025), AutoCodeRover (Zhang et al., 2024) and SWE-agent (Yang et al., 2024). RepairAgent (Bouzenia et al., 2025) fixes bugs based on test cases, and guides the agent by defining a finite-state machine that outlines the typical steps followed by a developer. Constraining the agent’s autonomy in this way can increase trust by aligning the agent with the typical workflow followed by a human developer. AutoCodeRover, a spinoff acquired by SonarSource, can solve natural language issues requiring bug fixing or feature addition. It establishes technical trust, e.g., by using program analysis on abstract syntax trees, and human trust, e.g., by extracting the intent of the software, which can be used to provide explanations of the suggested code edits. SWE-agent (Yang et al., 2024) follows a philosophy similar to Devin, by making file navigation tools and interfaces available to an AI software engineer. It works on a text-based program representation and does not employ any program analysis.
@@ 4. Establishing Trust
What makes us trust human-written code, but not necessarily the code generated by an LLM? Part of the reason is the perceived capability of “passing the blame”. If a human developer is involved, there is the promise of getting feedback from the developer as needed. Of course, this does not always hold, e.g., if the developer eventually leaves an organization. Nevertheless, accepting a code commit from a developer often partially depends on the reputation of the developer within the organization. For an AI software engineer in the form of an LLM agent to earn a reputation similar to a senior human colleague, it has to integrate established quality assurance techniques and collaborate effectively with human developers. We outline several ideas toward this goal.
Testing and Lightweight Static Analysis
One way to increase technical trust is to retrofit testing and lightweight static analyzers into an LLM agent. For example, in the process of code generation, additional artifacts, such as tests that exercise the newly added code, can be generated as well (Ryan et al., 2024), possibly derived from a natural language description of the agent’s task. A crucial challenge is to create appropriate test oracles that check the actual outputs against expected outputs, e.g., by inferring the expected outputs from natural language specifications.
@@
An enhanced degree of technical trust can come from formal proofs. A promising paradigm in this regard is automated, proof-oriented programming (Chakraborty et al., 2025). In this paradigm, LLMs generate the code together with the necessary formal specifications (pre/post-conditions, loop invariants, and so on) in a verifiable language, such as F*, Dafny, or Verus. Such programs can then be automatically verified, providing greater trust than testing alone.
Guardrails for Increased Security and Alignment
Trust can also be ensured through the use of guardrails.These can serve as a sanitization mechanisms, filtering malicious inputs before they reach the LLM and validating the generated code before it is returned to a user. Guardrails help defend against three key threats: prompt injections, where prompts trick the LLM into bypassing safety measures; malicious code, where harmful input code leads the LLM to generate or modify malicious code; and vulnerable code, where even unintentionally insecure input may cause the LLM to propagate or alter undesired code.
Specification Inference for Explainability
A more conceptual mechanism to establish both technical and human trust would be to infer the code intent from the initial, possibly buggy program. The system-level intent of what a large software system is supposed to do can often be crisply captured by a detailed natural language prompt. What is missing is the intent of the functions or methods. An LLM agent could be geared towards such specification inference, navigating the code base via code search, and trying to infer the intended behavior (Ruan et al., 2025). Such explicit unit-level specification inference can enable the program modifications to be accompanied by justifications.
Effective AI-Human Collaboration
A key factor for increasing human trust into AI software engineers will be to enforce effective AI-human collaboration patterns. As experienced first-hand by the AutoCodeRover team (Zhang et al., 2024) via real-life anecdotes communicated by clients, developer hesitation in accepting AI-generated code also comes from the volume of code that can be quickly generated by AI tools, overwhelming human developers. Supporting AI generated code with confidence scores can help reduce developer hesitation. Moreover, personalizing LLM agents to the experiences and expertise of specific users can help in increasing familiarity with AI-suggested code. The latest release of the Codex agent (OpenAI, 2025) in May 2025 (the Codex agent can run synchronously in the terminal watching a human, while also running long-running tasks in the cloud asynchronously) sharpens these questions – can a LLM coding agent be trusted as an intern or a junior team member?
Realistic Expectations and Critical Assessment
Clear documentation of an agent’s capabilities and known limitations can support informed decision making and can set realistic expectations. Any social biases being implicitly introduced by agents will need to be assessed. Integrating such assessments into the day-to-day development workflows enables teams to continuously monitor and improve the trustworthiness.
@@ 5. Outlook
As AI software engineers take on more core development tasks, their success will hinge not just on technical capability but on earning developer trust. Rather than relying on many separate agents for specialized software engineering tasks, it could be worthwhile to create a unified software engineering agent that combines coding, testing, debugging, etc. into a coherent, explainable workflow. To become a trusted collaborator, such an agent must offer transparency, adapt to feedback, and integrate safeguards that ensure quality and security. Ultimately, programming with AI will mean not full automation, but effective delegation – where human oversight and AI support work hand in hand.
@@ Acknowledgements.
- OpenAI (2025) OpenAI. 2025. Codex and Codex CLI. https://openai.com/codex/.